# Rigorous dense graph limit of a model for

### v1

@@ 4. Reformulation of the model and main results
Fix $\gamma>1$, $\nu>0$, $\lambda>0$, and $r>0$, and let the assumptions (3.1)–(3.6) hold. Let $B^{N}\in\mathbb{B}^{N}_{r}$, $N\in\mathbb{N}$, be a sequence of global minimizers of the discrete energy functionals $F^{N}$, given by (4.1)–(4.2), on the sets $\mathbb{B}^{N}_{r}$. Then $b^{N}=\mathcal{Q}^{N}[B^{N}]$ converges weakly in $L^{2}(\mathcal{I}\times\mathcal{I})$ to $b\in L^{2}(\mathcal{I}\times\mathcal{I})$, which is the global minimizer of the continuum energy functional $\mathcal{F}$, given by (4.4)–(4.5), on $L_{r}^{\omega}(\mathcal{I}\times\mathcal{I})$ with $\omega=\max\{2,\gamma\}$.
@@ 5. The Kirchhoff law and the Poisson equation
$\sqrt{2}\left\|p\right\|_{L^{2}(\mathcal{I})}=\left\|\mathfrak{d}p\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}.$ (5.1)
@@
$\left\|\mathfrak{d}p\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}^{2}$ $=$ $\int_{\mathcal{I}}\int_{\mathcal{I}}(p(x)-p(y))^{2}\,\mathrm{d}x\,\mathrm{d}y$
@@
$\int_{\mathcal{I}}\int_{\mathcal{I}}(z(x)-z(y))^{2}\,\mathrm{d}w(x,y)\geq\lambda\int_{\mathcal{I}}z(x)^{2}\,\mathrm{d}x\qquad\mbox{for all }z\in L^{2}_{0}(\mathcal{I}).$ (5.3)
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\mathfrak{d}p(x,y))(\mathfrak{d}\varphi(x,y))\,\mathrm{d}w^{N}(x,y)=\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}(\varphi_{i}-\varphi_{j})W_{ij}^{N}B_{ij}\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\mathfrak{d}p(x,y)\,\mathrm{d}x\,\mathrm{d}y,$
@@
$\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\mathfrak{d}p(x,y)\,\mathrm{d}x\,\mathrm{d}y=\frac{1}{N}\int_{\mathcal{I}_{i}^{N}}p(x)\,\mathrm{d}x-\frac{1}{N}\int_{\mathcal{I}_{j}^{N}}p(y)\,\mathrm{d}y=\frac{1}{N^{2}}\left(\mathcal{Z}^{N}[p]_{i}-\mathcal{Z}^{N}[p]_{j}\right),$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\mathfrak{d}p^{N}(x,y))(\mathfrak{d}\varphi(x,y))\,\mathrm{d}w^{N}(x,y)=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}(\varphi_{i}-\varphi_{j})(p^{N}_{i}-p^{N}_{j})W_{ij}^{N}B_{ij},$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\mathfrak{d}p)^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x<\infty$ (5.6)
@@
$\lambda r\left\|p^{n}\right\|_{L^{2}(\mathcal{I})}^{2}\leq\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}|\mathfrak{d}p^{n}|^{2}\,\mathrm{d}w(x,y)$ $=$ $2\int_{\mathcal{I}}\sigma p^{n}\,\mathrm{d}x$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\mathfrak{d}p(x,y)\mathfrak{d}\varphi(x,y)\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi(x)\,\mathrm{d}x$ (5.10)
for all test functions $\varphi\in L_{0}^{\infty}(\mathcal{I})$. Noting that (5) also implies a uniform bound on $\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}|\mathfrak{d}p^{n}|^{2}\,\mathrm{d}w(x,y)$, we have due to the weak lower semicontinuity of the $L^{2}$-norm,
$\int_{\mathcal{I}}\int_{\mathcal{I}}b(\mathfrak{d}p)^{2}\,\mathrm{d}w(x,y)=\left\|\sqrt{bw}(\mathfrak{d}p)\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}^{2}\leq\liminf_{n\to\infty}\left\|\sqrt{b^{n}w}(\mathfrak{d}p^{n})\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}^{2}<+\infty,$
where for the identification of the limit we used the fact that $\sqrt{b^{n}}\to\sqrt{b}$ in the norm topology of $L^{4}(\mathcal{I}\times\mathcal{I})$, recalling that $b^{n}\geq r>0$, and $\mathfrak{d}p^{n}\rightharpoonup\mathfrak{d}p$ weakly in $L^{2}(\mathcal{I}\times\mathcal{I})$. Consequently, we can use a sequence of cut-off versions of $p$ as test functions in (5.10) and pass to the limit to remove the cut-off. This gives
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\mathfrak{d}p)^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x,$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\mathfrak{d}p^{N})^{2}\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma p^{N}\,\mathrm{d}x<\infty,$ (5.11)
@@ 6. The energy functionals
$\frac{\nu}{\gamma}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)^{\gamma}\left[\ell^{N}(x,y)\right]^{\gamma+1}\,\mathrm{d}w^{N}(x,y)$ $=$ $\frac{\nu}{\gamma}\sum_{i=1}^{N}\sum_{j=1}^{N}W_{ij}^{N}B_{ij}^{\gamma}\left(L_{ij}^{N}\right)^{\gamma+1}\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\,\mathrm{d}x\,\mathrm{d}y$
$=$ $\frac{\nu}{\gamma N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}W_{ij}^{N}B_{ij}^{\gamma}\left(L_{ij}^{N}\right)^{\gamma+1},$
where we used the fact that, by construction, $b^{N}=\mathcal{Q}^{N}[B]\equiv B_{ij}$ on the patch $\mathcal{I}_{i}^{N}\times\mathcal{I}_{j}^{N}$, and similarly $\ell^{N}=\mathcal{Q}^{N}[L^{N}]\equiv L^{N}_{ij}$, $w^{N}=\mathcal{Q}^{N}[W^{N}]\equiv W^{N}_{ij}$ on $\mathcal{I}_{i}^{N}\times\mathcal{I}_{j}^{N}$.
@@
$\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}p^{N}(x,y)\right)^{2}\,\mathrm{d}w^{N}(x,y)=\sum_{i=1}^{N}\sum_{j=1}^{N}W_{ij}^{N}B_{ij}\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\left(\mathfrak{d}p^{N}(x,y)\right)^{2}\,\mathrm{d}x\,\mathrm{d}y,$
@@
$\mathcal{F}^{N}[b]=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left(B_{ij}(P_{i}-P_{j})^{2}+\frac{\nu}{\gamma}B_{ij}^{\gamma}\left(L_{ij}^{N}\right)^{\gamma+1}\right)W_{ij}^{N}=F^{N}[B].$
@@
$\mathcal{F}_{\mathrm{kin}}[b]:=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b\left(\mathfrak{d}p\right)^{2}\,\mathrm{d}w(x,y),$ (6.1)
@@
$\mathcal{F}_{\mathrm{kin}}^{N}[b]:=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b\left(\mathfrak{d}p^{N}\right)^{2}\,\mathrm{d}w^{N}(x,y),$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\mathfrak{d}p^{N}(x,y))(\mathfrak{d}\varphi^{N}(x,y))\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi^{N}(x)\,\mathrm{d}x$ (6.2)
@@
$(\mathfrak{d}p^{N})(\mathfrak{d}\varphi^{N})w^{N}=\mathcal{Z}^{N}\left[(\mathfrak{d}p^{N})(\mathfrak{d}\varphi^{N})w^{N}\right].$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b](x,y)(\mathfrak{d}p^{N}(x,y))(\mathfrak{d}\varphi^{N}(x,y))\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi^{N}(x)\,\mathrm{d}x.$
@@
$\mathcal{F}_{\mathrm{kin}}^{N}[b]$ $=$ $\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\mathfrak{d}p^{N}(x,y))^{2}\,\mathrm{d}w^{N}(x,y)$
@@
$=$ $\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b](x,y)(\mathfrak{d}p^{N}(x,y))^{2}\,\mathrm{d}w^{N}(x,y)$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}|\mathfrak{d}p^{n}|^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p^{n}\,\mathrm{d}x\,\mathrm{d}y.$ (6.3)
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\mathfrak{d}p)^{2}\,\mathrm{d}x\,\mathrm{d}y=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x.$
@@
$\lim_{n\to\infty}\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}(\mathfrak{d}p^{n})^{2}\,\mathrm{d}x\,\mathrm{d}y=\lim_{n\to\infty}\int_{\mathcal{I}}\sigma p^{n}\,\mathrm{d}x=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\mathfrak{d}p)^{2}\,\mathrm{d}x\,\mathrm{d}y,$
@@
Our next result shows that the functional $\mathcal{F}_{\mathrm{kin}}$ is convex on its domain. The proof goes along the lines of [11, Proposition 3.2] or [12, Section 3].
@@
Finiteness of of $\mathcal{F}$ on $L^{2}_{r}(\mathcal{I}\times\mathcal{I})$ follows directly from (5.6). To prove convexity, we use the identity (5.6),
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\mathfrak{d}p)^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x,$
where $p=p[b]\in L^{2}_{0}(\mathcal{I})$ the unique solution of the Poisson equation (4.5). We calculate the second-order variation of the term $\int_{\mathcal{I}}\sigma p\,\mathrm{d}x$. Let us denote $\frac{\delta p[b;\Phi]}{\delta b}$ the first variation of $p=p[b]$ in direction $\Phi\in L^{2}_{+}(\mathcal{I}\times\mathcal{I})$. Using this as a test function in the weak formulation (4.5) and calculating the first-order variation of the resulting expression in the same direction gives

### v2

@@ 4. Reformulation of the model and main results
Fix $\gamma>2$, $\nu>0$, $\lambda>0$, and $r>0$, and let the assumptions (3.1)–(3.6) hold. Let $B^{N}\in\mathbb{B}^{N}_{r}$, $N\in\mathbb{N}$, be a sequence of global minimizers of the discrete energy functionals $F^{N}$, given by (4.1)–(4.2), on the sets $\mathbb{B}^{N}_{r}$. Then $b^{N}=\mathcal{Q}^{N}[B^{N}]$ converges weakly in $L^{2}(\mathcal{I}\times\mathcal{I})$ to $b\in L^{2}(\mathcal{I}\times\mathcal{I})$, which is the global minimizer of the continuum energy functional $\mathcal{F}$, given by (4.4)–(4.5), on $L_{r}^{\gamma}(\mathcal{I}\times\mathcal{I})$.
@@ 5. The Kirchhoff law and the Poisson equation
$\sqrt{2}\left\|p\right\|_{L^{2}(\mathcal{I})}=\left\|\overline{\nabla}p\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}.$ (5.1)
@@
$\left\|\overline{\nabla}p\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}^{2}$ $=$ $\int_{\mathcal{I}}\int_{\mathcal{I}}(p(x)-p(y))^{2}\,\mathrm{d}x\,\mathrm{d}y$
@@
$\int_{\mathcal{I}}\int_{\mathcal{I}}\big(\overline{\nabla}z(x,y)\big)^{2}\,\mathrm{d}w(x,y)\geq\lambda\int_{\mathcal{I}}z(x)^{2}\,\mathrm{d}x\qquad\mbox{for all }z\in L^{2}_{0}(\mathcal{I}).$ (5.3)
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)$ $(\overline{\nabla}p(x,y))(\overline{\nabla}\varphi(x,y))\,\mathrm{d}w^{N}(x,y)$
$=\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}(\varphi_{j}-\varphi_{i})W_{ij}^{N}B_{ij}\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\overline{\nabla}p(x,y)\,\mathrm{d}x\,\mathrm{d}y,$
@@
$\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\overline{\nabla}p(x,y)\,\mathrm{d}x\,\mathrm{d}y=\frac{1}{N}\int_{\mathcal{I}_{j}^{N}}p(y)\,\mathrm{d}y-\frac{1}{N}\int_{\mathcal{I}_{i}^{N}}p(x)\,\mathrm{d}x=\frac{1}{N^{2}}\left(\mathcal{Z}^{N}[p]_{j}-\mathcal{Z}^{N}[p]_{i}\right),$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)$ $(\overline{\nabla}p^{N}(x,y))(\overline{\nabla}\varphi(x,y))\,\mathrm{d}w^{N}(x,y)$
$=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}(\varphi_{i}-\varphi_{j})(p^{N}_{i}-p^{N}_{j})W_{ij}^{N}B_{ij},$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\overline{\nabla}p)^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x<\infty$ (5.6)
@@
$\lambda r\left\|p^{n}\right\|_{L^{2}(\mathcal{I})}^{2}\leq\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}|\overline{\nabla}p^{n}|^{2}\,\mathrm{d}w(x,y)$ $=$ $2\int_{\mathcal{I}}\sigma p^{n}\,\mathrm{d}x$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\overline{\nabla}p(x,y)\overline{\nabla}\varphi(x,y)\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi(x)\,\mathrm{d}x$ (5.10)
for all test functions $\varphi\in L_{0}^{\infty}(\mathcal{I})$. Noting that (5) also implies a uniform bound on $\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}|\overline{\nabla}p^{n}|^{2}\,\mathrm{d}w(x,y)$, we have due to the weak lower semicontinuity of the $L^{2}$-norm,
$\int_{\mathcal{I}}\int_{\mathcal{I}}b(\overline{\nabla}p)^{2}\,\mathrm{d}w(x,y)=\left\|\sqrt{bw}(\overline{\nabla}p)\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}^{2}\leq\liminf_{n\to\infty}\left\|\sqrt{b^{n}w}(\overline{\nabla}p^{n})\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}^{2}<\infty,$
where for the identification of the limit we used the fact that $\sqrt{b^{n}}\to\sqrt{b}$ in the norm topology of $L^{4}(\mathcal{I}\times\mathcal{I})$, recalling that $b^{n}\geq r>0$, and $\overline{\nabla}p^{n}\rightharpoonup\overline{\nabla}p$ weakly in $L^{2}(\mathcal{I}\times\mathcal{I})$. Consequently, we can use a sequence of cut-off versions of $p$ as test functions in (5.10) and pass to the limit to remove the cut-off. This gives
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\overline{\nabla}p)^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x,$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\overline{\nabla}p^{N})^{2}\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma p^{N}\,\mathrm{d}x<\infty,$ (5.11)
@@ 6. The energy functionals
$\frac{\nu}{\gamma}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)^{\gamma}\left[\ell^{N}(x,y)\right]^{\gamma+1}\,\mathrm{d}w^{N}(x,y)$ $=$ $\frac{\nu}{\gamma}\sum_{i=1}^{N}\sum_{j=1}^{N}W_{ij}^{N}B_{ij}^{\gamma}\left(\mathrm{L}_{ij}^{N}\right)^{\gamma+1}\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\,\mathrm{d}x\,\mathrm{d}y$
$=$ $\frac{\nu}{\gamma N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}W_{ij}^{N}B_{ij}^{\gamma}\left(\mathrm{L}_{ij}^{N}\right)^{\gamma+1},$
where we used the fact that, by construction, $b^{N}=\mathcal{Q}^{N}[B]\equiv B_{ij}$ on the patch $\mathcal{I}_{i}^{N}\times\mathcal{I}_{j}^{N}$, and similarly $\ell^{N}=\mathcal{Q}^{N}[\mathrm{L}^{N}]\equiv\mathrm{L}^{N}_{ij}$, $w^{N}=\mathcal{Q}^{N}[W^{N}]\equiv W^{N}_{ij}$ on $\mathcal{I}_{i}^{N}\times\mathcal{I}_{j}^{N}$.
@@
$\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\overline{\nabla}p^{N}(x,y)\right)^{2}\,\mathrm{d}w^{N}(x,y)=\sum_{i=1}^{N}\sum_{j=1}^{N}W_{ij}^{N}B_{ij}\int_{\mathcal{I}_{i}^{N}}\int_{\mathcal{I}_{j}^{N}}\left(\overline{\nabla}p^{N}(x,y)\right)^{2}\,\mathrm{d}x\,\mathrm{d}y,$
@@
$\mathcal{F}^{N}[b]=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left(B_{ij}(P_{i}-P_{j})^{2}+\frac{\nu}{\gamma}B_{ij}^{\gamma}\left(\mathrm{L}_{ij}^{N}\right)^{\gamma+1}\right)W_{ij}^{N}=F^{N}[B].$
@@
$\mathcal{F}_{\mathrm{kin}}[b]:=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b\left(\overline{\nabla}p\right)^{2}\,\mathrm{d}w(x,y),$ (6.1)
@@
$\mathcal{F}_{\mathrm{kin}}^{N}[b]:=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b\left(\overline{\nabla}p^{N}\right)^{2}\,\mathrm{d}w^{N}(x,y),$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\overline{\nabla}p^{N}(x,y))(\overline{\nabla}\varphi^{N}(x,y))\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi^{N}(x)\,\mathrm{d}x$ (6.2)
@@
$(\overline{\nabla}p^{N})(\overline{\nabla}\varphi^{N})w^{N}=\mathcal{Z}^{N}\left[(\overline{\nabla}p^{N})(\overline{\nabla}\varphi^{N})w^{N}\right].$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b](x,y)(\overline{\nabla}p^{N}(x,y))(\overline{\nabla}\varphi^{N}(x,y))\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi^{N}(x)\,\mathrm{d}x.$
@@
$\mathcal{F}_{\mathrm{kin}}^{N}[b]$ $=$ $\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\overline{\nabla}p^{N}(x,y))^{2}\,\mathrm{d}w^{N}(x,y)$
@@
$=$ $\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b](x,y)(\overline{\nabla}p^{N}(x,y))^{2}\,\mathrm{d}w^{N}(x,y)$
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}|\overline{\nabla}p^{n}|^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p^{n}\,\mathrm{d}x\,\mathrm{d}y.$ (6.3)
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\overline{\nabla}p)^{2}\,\mathrm{d}x\,\mathrm{d}y=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x.$
@@
$\lim_{n\to\infty}\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b^{n}(\overline{\nabla}p^{n})^{2}\,\mathrm{d}x\,\mathrm{d}y=\lim_{n\to\infty}\int_{\mathcal{I}}\sigma p^{n}\,\mathrm{d}x=\int_{\mathcal{I}}\sigma p\,\mathrm{d}x=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(\overline{\nabla}p)^{2}\,\mathrm{d}x\,\mathrm{d}y,$
@@
Our next result shows that the functional $\mathcal{F}_{\mathrm{kin}}$ is convex on its domain. Convexity of the kinetic part of the discrete energy functional (1.4) and its continuum counterpart has been established in [11, Proposition 3.2] and in [12, Section 3], based on a proof of positive semidefinitness of the Hessian. Here we offer a rigorous proof using the variational formulation of the Poisson equation (4.5).
@@
Finiteness of $\mathcal{F}_{\mathrm{kin}}$ on $L^{2}_{r}(\mathcal{I}\times\mathcal{I})$ follows directly from (5.6). To prove convexity on $L^{2}_{r}(\mathcal{I}\times\mathcal{I})$, we first prove convexity on $L_{r}^{\infty}(\mathcal{I}\times\mathcal{I})$. According to Lemma 5.4, for $b\in L_{r}^{\infty}(\mathcal{I}\times\mathcal{I})$, the Poisson equation (4.5) is well posed in $L^{2}_{0}(\mathcal{I})$ with test functions in $L^{2}_{0}(\mathcal{I})$. Moreover, its unique solution $p=p[b]\in L^{2}_{0}(\mathcal{I})$ is the unique minimizer of the functional
$\mathcal{J}[v;b]:=\frac{1}{4}\iint_{\mathcal{I}\times\mathcal{I}}b(x,y)\,(\overline{\nabla}v(x,y))^{2}\,\,\mathrm{d}w(x,y)-\int_{\mathcal{I}}\sigma(x)\,v(x)\,\,\mathrm{d}x\qquad\text{over }v\in L^{2}_{0}(\mathcal{I}).$
Indeed, (4.5) is the Euler–Lagrange equation of $\mathcal{J}[\cdot;b]$. Testing (4.5) with $\varphi=p$ yields $\frac{1}{2}\iint_{\mathcal{I}\times\mathcal{I}}b(\overline{\nabla}p)^{2}\,\,\mathrm{d}w=\int_{\mathcal{I}}\sigma p\,\,\mathrm{d}x$, hence
@@
$\mathcal{J}[p;b]=\frac{1}{4}\iint_{\mathcal{I}\times\mathcal{I}}b(\overline{\nabla}p)^{2}\,\,\mathrm{d}w-\int_{\mathcal{I}}\sigma p\,\,\mathrm{d}x=\frac{1}{2}\,\mathcal{F}_{\mathrm{kin}}[b]-\mathcal{F}_{\mathrm{kin}}[b]=-\frac{1}{2}\,\mathcal{F}_{\mathrm{kin}}[b],$
and, consequently,